OpenStudy (anonymous):

Simplify. (3a – 4)^2 A. 9a^2 – 16 B. 9a^2 – 24a + 16 C. 6a^2– 24a + 8 D. 9^2a – 7 + 16

OpenStudy (paki):

simple... just use the formula.... (a-b)^2=a^2-2(a)(b)+b^2 to solve.....

OpenStudy (anonymous):

uhhh i have no idea how to do that

OpenStudy (paki):

ok see... here at (3a – 4)^2.... we have a=3a and b= -4... now put these in the formula... and solve...

OpenStudy (anonymous):

okay hold on ill try

OpenStudy (anonymous):

yeah i still dont get it at all :/

OpenStudy (paki):

hmmm... ok see... this is the formula... (a-b)^2=a^2-2(a)(b)+b^2 (3a – 4)^2 = a=3a and b=4... put in the formula... (3a – 4)^2 = (3a)^2-2(3a)(4)+(4)^2 (3a – 4)^2 = 9a^2-24a+16... this is the solution... got it now...

OpenStudy (anonymous):

i still dont get it but ill try to study the formula until i can solve it on my own i guess